OpenStudy (bimbim2697):

Use integer coffefficient,What is the standards form of the equations below (y-3)=2/3(x+6)

OpenStudy (anonymous):

y - 3 = 2x/3 + 4 y = 2x/3 + 7 2x/3 - y = -7

OpenStudy (bimbim2697):

WHAT IF I HAD 2X+3Y=7 WOULD THT BE WRONG?

OpenStudy (anonymous):

Yes. But, you can re-write my equation by multiplying by 3 throughout! 2x - 3y = -21

OpenStudy (bimbim2697):

OH THANKS MUCH U SEE MY OTHER QUESTIONS BELOW